# KKL Observer Synthesis for Nonlinear Systems via Physics-Informed Learning

### v1

@@ 5.2 Technical Lemmas
The lemmas below, though familiar in various forms, are presented here with precise quantification of parameters, which are essential for deriving input-to-state stability bounds of the estimation error in the following sections. Moreover, these bounds are also used in Section 8 to explicitly compute the thresholds for detecting and isolating sensor faults.
Recall $z=\mathcal{T}(x)$, where
$\dot{z}(t)=Az(t)+Bh(\bar{x}(t))$
with $\bar{x}(t)\coloneqq x(t;x_{0},0)$ the noise-free state trajectory. Also, recall the dynamics of $\hat{z}(t)=\hat{\mathcal{T}}_{\theta}(\hat{x}(t))$ given in (19).
@@
Suppose the matrix $A$ is Hurwitz and diagonalizable with eigenvalue decomposition $A=V\Lambda V^{-1}$. Then,
$\|\exp(At)\|\leq\text{cond}(V)e^{\lambda_{\min}(A)t}$
@@
[…]
where $\text{cond}(V)=\|V\|\|V^{-1}\|$ is the condition number of $V$.
The proof of this lemma is straightforward. A similar result can be obtained when $A$ is not diagonalizable by using the Jordan form of $A$; see [51, Appendix C.5]. However, since diagonalizable matrices are dense in the space of square matrices and also since $A$ is a design matrix, the diagonalizability assumption is mild.
@@
$\|\tilde{z}(t)\|\leq\|\tilde{z}_{0}\|\text{cond}(V)e^{\lambda_{\min}(A)t}\\ +\frac{\text{cond}(V)}{|\lambda_{\min}(A)|}\|B\|(1-e^{\lambda_{\min}(A)t})\big{[}\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v}\big{]}$ (26)
@@
Define $\sigma(t)\coloneq h(\bar{x}(t))-y(t)=h(\bar{x}(t))-h(x(t))-v(t)$, where $x(t)\coloneqq x(t;x_{0},w)$ is the solution of (20). Then, from the dynamics $\dot{\tilde{z}}(t)=A\tilde{z}(t)+B\sigma(t)$, we have that the solution $\tilde{z}(t;\tilde{z}_{0},\sigma)$ satisfies
$\|\tilde{z}(t)\|\leq\|\exp(At)\tilde{z}_{0}\|+\int_{0}^{t}\|\exp(A\tau)B\sigma(t-\tau)\|\text{d}\tau.$ (27)
@@
[…]
$\hskip 56.9055pt\leq\int_{0}^{t}\|\exp(A\tau)B\|\|\sigma(t-\tau)\|\text{d}\tau$
$\hskip 56.9055pt\leq\int_{0}^{t}\|\exp(A\tau)B\|\text{d}\tau\,\max_{\tau\geq 0}\|\sigma(\tau)\|.$ (28)
From (5.1), it follows that
$\max_{\tau\geq 0}\|\sigma(\tau)\|\leq\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v}.$
The result follows by substituting the above inequality in (28), then in (27), and using Lemma 11.
@@ 5.3 Robustness to the Approximation Error
Given that the activation functions used in the neural network $\hat{\mathcal{T}}_{\eta}^{*}$ are Lipschitz continuous, e.g., Rectified Linear Unit (ReLU) networks, we show that the KKL observer (7) for the noise-free system (1) is robust to the neural network approximation error. We base our result on the following lemma.
Lemma 13.
Given a neural network $\hat{\mathcal{T}}_{\eta}^{*}$ with $l$ layers and, at each layer $i$, the weight matrix and the activation function are denoted by $W_{\eta}^{(i)}$ and $\sigma^{(i)}:\mathbb{R}\to\mathbb{R}$, respectively. Then, if $\sigma^{(i)}$ is Lipschitz continuous with constant $\ell_{\eta}^{i}$, for every $i\in\{1,\dots,l\}$, it holds that $\hat{\mathcal{T}}^{*}$ is Lipschitz continuous with constant $\ell_{\eta}=\ell_{\eta}^{1}\dots\ell_{\eta}^{l}\|W_{\eta}^{(1)}\|\dots\|W_{\eta}^{(l)}\|$, i.e., $\forall$ $\hat{z},z\in\mathbb{R}^{n_{z}}$,
$\|\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))-\hat{\mathcal{T}}_{\eta}^{*}(z(t))\|\leq\ell_{\eta}\|\hat{z}(t)-z(t)\|.$ (29)
@@
Let $\Sigma^{(i)}(b^{(i)})=[\begin{array}[]{ccc}\sigma^{(i)}(b_{1}^{(i)})&\dots&\sigma^{(i)}(b_{m^{(i)}}^{(i)})\end{array}]^{\top}$ be the output at layer $i$, where $b^{(i)}=W^{(i)}a^{(i-1)}\in\mathbb{R}^{m^{(i)}}$ with $m^{(i)}$ the number of neurons at layer $i$, and $a^{(i-1)}$ the input $z\in\mathbb{R}^{n_{z}}$ if $i=1$ or the activated output of layer $i-1$ if $i\geq 2$. Then, for every $\hat{a}^{(i-1)},a^{(i-1)}$,
$\begin{array}[]{l}\|\Sigma^{(i)}(W^{(i)}\hat{a}^{(i-1)})-\Sigma^{(i)}(W^{(i)}a^{(i-1)})\|\\ \qquad\qquad\qquad\leq\ell_{\eta}^{i}\|W^{(i)}\hat{a}^{(i-1)}-W^{(i)}a^{(i-1)}\|\\ \qquad\qquad\qquad\leq\ell_{\eta}^{i}\|W^{(i)}\|\|\hat{a}^{(i-1)}-a^{(i-1)}\|.\end{array}$
The result follows by applying the above inequality at each layer.
Remark 14.
Consider the ReLU activation function given by $\Sigma(b)=[\begin{array}[]{ccc}\max(0,b_{1})&\dots&\max(0,b_{m})\end{array}]^{\top}$, which is Lipschitz continuous with a Lipschitz constant equal to $1$ as $\|\Sigma(\hat{b})-\Sigma(b)\|\leq\|\hat{b}-b\|,$ where $b\in\mathbb{R}^{m}$. Thus, for ReLU networks, the Lipschitz constant $\ell_{\eta}$ is equal to the product of the maximum singular values of the weight matrices. Moreover, the weights $W_{\eta}^{(i)}$ for layer $i\in[l]$ can be regularized while training $\hat{\mathcal{T}}_{\eta}^{*}$ to restrict their singular values. Nevertheless, the Lipschitz constant given by (29) could be conservative and may result in very large values if the number of layers is large. Although it is an NP-hard problem [55], estimating a tighter bound of the Lipschitz constant of neural networks is a topic of interest in the machine learning community [53, 55, 22, 27]. If one is interested in the local Lipschitz continuity of a feedforward ReLU network, then [20] shows that the local Lipschitz constant estimation problem can be reduced to a semi-definite programming problem. $\diamond$
Assumption 15.
The neural networks $\hat{\mathcal{T}}_{\theta}$ and $\hat{\mathcal{T}}_{\eta}^{*}$ are Lipschitz continuous over $\mathcal{X}$ and $\mathcal{Z}$, respectively, where $\mathcal{Z}\supseteq\mathcal{T}(\mathcal{X})$. In particular, for every $x,\hat{x}\in\mathcal{X}$, there exists $\ell_{\theta}\in\mathbb{R}_{>0}$ such that
$\|\hat{\mathcal{T}}_{\theta}(x)-\hat{\mathcal{T}}_{\theta}(\hat{x})\|\leq\ell_{\theta}\|x-\hat{x}\|$
and, for every $z,\hat{z}\in\mathcal{Z}$, there exists $\ell_{\eta}\in\mathbb{R}_{>0}$ such that
$\|\hat{\mathcal{T}}_{\eta}^{*}(z)-\hat{\mathcal{T}}_{\eta}^{*}(\hat{z})\|\leq\ell_{\eta}\|z-\hat{z}\|.$
There exists a class $\mathcal{K}_{\infty}$ function $\alpha$ such that, for every $z\in\mathcal{Z}$, the neural network approximation error $\mathcal{E}^{*}(z)$ appearing in (18) satisfies
$\|\mathcal{E}^{*}(z)\|\leq\alpha(\|z\|).$
This is because $\|\mathcal{E}^{*}(z)\|\leq\|\mathcal{T}^{*}(z)\|+\|\hat{\mathcal{T}}_{\eta}^{*}(z)\|$. Then, the above inequality follows from (8) and (29) with $\alpha(\|z\|)=\rho(\|z\|)+\ell_{\eta}\|z\|$.
Because the state space $\mathcal{X}\subset\mathbb{R}^{n_{x}}$ is bounded, $h(\cdot)$ is a smooth map, and $A$ is Hurwitz, there exists a compact set $\mathcal{Z}\subset\mathbb{R}^{n_{z}}$ containing the trajectory $z(t;\mathcal{T}(x_{0}))$ of (5) for every $t\in\mathbb{R}_{\geq 0}$ and every $x_{0}\in\mathcal{X}$. Thus, as a consequence of (8) and (29), there exists a finite approximation bound $\epsilon^{*}>0$ satisfying
$\epsilon^{*}=\sup_{z\in\mathcal{Z}}\|\mathcal{E}^{*}(z)\|.$ (30)
There have been several attempts [50, 28, 26, 14] to estimate $\epsilon^{*}$ and show that it can be reduced by improving the design and learning technique of the neural network, and also by increasing the size (number of initialization points $p$ and sampling frequency $\tau/T$) of the dataset $(X,Z)$ (see [40]).
Proposition 16.
@@
Let Assumptions 2 and 6 hold. Suppose $A$ is diagonalizable with eigendecomposition $A=V\Lambda V^{-1}$. Let $x(t)$ be the state of the noise-free system (1) and $\hat{x}(t)$ be the estimate provided by (19). Then, for every $t\in\mathbb{R}_{\geq 0}$, the estimation error $\xi(t)=x(t)-\hat{x}(t)$ satisfies
$\|\xi(t)\|\leq ae^{\lambda_{\min}(A)t}+\epsilon^{*}$ (31)
where $\epsilon^{*}$ is given in (30) and $a=\ell_{\eta}\|\tilde{z}_{0}\|\text{cond}(V)$ with $\tilde{z}_{0}=z(t)-\hat{z}(0)$ the initialization error in the $z$-coordinate.
@@
We have
$\|\xi(t)\|$ $=\|\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))-\mathcal{T}^{*}(z(t))\|$
$=\|\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))-\hat{\mathcal{T}}_{\eta}^{*}(z(t))-\mathcal{E}^{*}(z(t))\|$
$\leq\|\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))-\hat{\mathcal{T}}_{\eta}^{*}(z(t))\|+\|\mathcal{E}^{*}(z(t))\|$
$\leq\ell_{\eta}\|\hat{z}(t)-z(t)\|+\epsilon^{*}$ (32)
where the first step is due to (18), the second step is due to the triangle inequality, and the last step is due to (29) and (30). Finally, using the inequality (26) with $\overline{w}=\overline{v}=0$ in (5.3), we obtain (31).
Inequality (31) says that the estimation error exponentially converges to a ball whose radius is dictated by a supremum of the approximation error of the inverse map. This indicates that one can achieve better estimation performance by improving the learning of $\hat{\mathcal{T}}_{\eta}^{*}$. A general rule of thumb for improving the learning of $\hat{\mathcal{T}}^{*}$ is to choose a deeper neural network and generate more training data through simulations.
Remark 17.
Notice that uniformly injective $\mathcal{T}$ implies the uniform injectivity of its inverse $\mathcal{T}^{*}$. It is important to note that uniform injectivity and Assumption 15 imply boundedness of approximation errors, i.e.,
$\|\mathcal{T}(x)-\hat{\mathcal{T}}_{\theta}(x)\|<\infty,\;\forall x\in\mathcal{X}$
and
$\|\mathcal{T}^{*}(z)-\hat{\mathcal{T}}_{\eta}^{*}(z)\|<\infty,\;\forall z\in\mathcal{Z}$
where $\mathbb{R}^{n_{z}}\supset\mathcal{Z}\supseteq\mathcal{T}(\mathcal{X})$. This holds because $\mathcal{X}\subset\mathbb{R}^{n_{x}}$ is a compact set and
$\|\mathcal{T}(x)-\hat{\mathcal{T}}_{\theta}(x)\|\leq\|\mathcal{T}(x)\|+\|\hat{\mathcal{T}}_{\theta}(x)\|\leq\varrho(\|x\|)+\ell_{\theta}\|x\|$
where $\varrho$ is given in (9) and $\ell_{\theta}$ is given in Assumption 15. Further improvement on this bound can be achieved practically by using the universal approximation property of neural networks, given that an ample amount of data is generated and an appropriate network architecture is chosen. $\diamond$
5.4 Robustness to Model Uncertainties and Sensor Noise
We now analyze the robustness of the learned observer (19) to estimate the state of an uncertain nonlinear system (20). Note that the design method of KKL observers as presented in Sections 3 and 4 remains the same for (20).
Proposition 18.
Let Assumptions 2, 6, 10, and 15 hold. Suppose $A$ is diagonalizable with eigendecomposition $A=V\Lambda V^{-1}$. Let $x(t)$ be the state of the uncertain system (20) and $\hat{x}(t)$ be the estimate provided by (19). Then, for every $t\in\mathbb{R}_{\geq 0}$, the estimation error $\xi(t)=x(t)-\hat{x}(t)$ satisfies
$\|\xi(t)\|\leq ae^{\lambda_{\min}(A)t}+b(t)[\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v}]+\epsilon^{*}$ (33)
where $\overline{w},\overline{v},\psi$ are given in Assumption 10, $\epsilon^{*}$ is given in (30), and
$a$ $=\ell_{\eta}\|\tilde{z}_{0}\|\text{cond}(V)$ (34a)
$b(t)$ $=\ell_{\eta}\frac{\text{cond}(V)}{|\lambda_{\min}(A)|}\|B\|(1-e^{\lambda_{\min}(A)t}).$ (34b)
{pf}
By substituting (26) into (5.3), we obtain (33).
Given that the model uncertainties and sensor noise are bounded, the above result shows that the KKL observer is robust in terms of input-to-state stability of the estimation error; see [52]. Moreover, it can be observed that improving the learning of $\hat{\mathcal{T}}_{\eta}^{*}$ while choosing $|\lambda_{\min}(A)|$ to be large and $\text{cond}(V)$ to be small [^6] For instance, when $A$ is a diagonal matrix, we have $V=I_{n_{z}}$ and $\text{cond}(V)=1$. reduces the estimation error.
6 Optimal KKL Observer Parameters
Choosing matrices $A$ and $B$ such that KKL observer (19) is robust to uncertainties is crucial. However, increasing robustness may inadvertently degrade learning accuracy and introduce substantial approximation errors. From (33), we observe that the stability of $A$ yields robustness, but it degrades controllability of the pair $(A,B)$, which is also a crucial requirement for KKL observers. To elucidate, more stable $A$ means more energy is required to steer the state $z(t)$ of (5) through the injected output $y(t)$. In this section, we present a design method for $A$ and $B$ that finds a trade-off between robustness and learnability of a KKL observer subject to the constraints that $A$ is Hurwitz and $(A,B)$ is controllable. To this end, we adopt techniques from $\mathcal{H}_{\infty}$-observer design and minimum energy control [19].
Recall the error dynamics (22) in the $z$-coordinates. To minimize the effect of $\sigma$ on the error $\tilde{z}$, we use the $\mathcal{H}_{\infty}$-based design of $A$ and $B$ under the constraint that $(A,B)$ is controllable. Consider the transfer function $G(s)=(sI-A)^{-1}B$ of (22). Then, one has
$\|G(s)\|_{\mathcal{H}_{\infty}}<\gamma,~{}\text{for some}~{}\gamma>0$
if and only if [19, Theorem 5.3] the bilinear matrix inequality (BMI)
$\left[\begin{array}[]{ccc}PA+A^{\top}P&PB&I_{n_{z}}\\ B^{\top}P&-\gamma I_{n_{y}}&0\\ I_{n_{z}}&0&-\gamma I_{n_{z}}\end{array}\right]<0$ (35)
is feasible for matrices $A\in\mathbb{R}^{n_{z}\times n_{z}}$ and $B\in\mathbb{R}^{n_{z}\times n_{y}}$, and a positive definite matrix $P=P^{\top}\in\mathbb{R}^{n_{z}\times n_{z}}$. Because $PA+A^{\top}P$ must be negative definite for (35) to hold, the resulting matrix $A$ will be Hurwitz by the Lyapunov theorem [47, Theorem 2.2.1].
Notice that merely solving (35) will yield a Hurwitz $A$ typically with very large eigenvalues, which may destroy the “practical” controllability from an energy point-of-view. That is, when the magnitude of the eigenvalues of $A$ is large, the required energy of the signal $y(t)$ to steer the state $z(t)$ from $z(0)=0$ to some $z(\tau)\in\mathbb{R}^{n_{z}}$, $\tau\in\mathbb{R}_{>0}$, would be impractically large. This will decrease the effect of the measured output $y(t)$ of the state estimation algorithm (7), leading to slow convergence to the true state. Thus, it is recommended to maximize the minimum eigenvalue $\lambda_{\min}(W)$ of the controllability gramian $W$ of (22), which is a worst-case controllability metric [43]. The controllability gramian $W=W^{\top}>0$ is the unique solution of the Lyapunov equation
$AW+WA^{\top}+BB^{\top}=0.$
We approximately rewrite this Lyapunov equation as a BMI
$-\varepsilon I\leq AW+WA^{\top}+BB^{\top}\leq\varepsilon I$ (36)
for a small $\varepsilon>0$. The minimum eigenvalue of $W$ is lower bounded by $\lambda>0$, i.e., $\lambda\leq\lambda_{\min}(W)$, if and only if [19, Lemma 1.1]
$\lambda I_{n_{z}}\leq W.$ (37)
The BMI feasibility problem for optimal $A$ and $B$ is then formulated as follows:
[…]
with respect to $\gamma>0$, $\lambda>0$, $A\in\mathbb{R}^{n_{z}\times n_{z}}$, $B\in\mathbb{R}^{n_{z}\times n_{y}}$, $P=P^{\top}>0$, $W=W^{\top}>0$, where
$g(\gamma,\lambda)\coloneqq c\gamma-(1-c)\lambda.$
Here, $\gamma$ is the $\mathcal{H}_{\infty}$ gain or a level of noise attenuation, $\lambda$ is the worst-case controllability metric, and $c\in[0,1]$ is a trade-off parameter. The choice of $c$ close to one indicates better robustness but slower observer transient while $c$ close to zero indicates better controllability but larger sensitivity to noise.
The problem (38) is not convex, quasi-convex, or even local-global [25]. However, there are several tools (e.g., PENLAB [23]) that can be used to obtain a locally optimal solution. Another strategy to solve (38) locally is to introduce certain relaxations such as iterative convex overbounding [56] and Young’s relation [58, 57], which convert the BMIs (35) and (36) into linear matrix inequalities (LMIs). Nonetheless, the simplest way to convert (38) into an LMI is to choose an appropriate positive definite matrix $P$ and controllability gramian $W$, then solve (38) to obtain $A$ and $B$. The matrices $A$ and $B$ obtained from (38) are guaranteed to satisfy the condition of Theorem 7 while also having good noise attenuation properties.

### v2

@@ 7.2 Technical Lemmas
The lemmas below are needed in proving the theorems later in this section. Recall $z=\mathcal{T}(x)$, where $\dot{z}(t)=Az(t)+Bh(\bar{x}(t))$ with $\bar{x}(t)\coloneqq x(t;x_{0})$ the noise-free state trajectory. Also, recall the dynamics of $\hat{z}(t)=\hat{\mathcal{T}}_{\theta}(\hat{x}(t))$ given in (26).
@@
Suppose $A$ is Hurwitz and diagonalizable with eigenvalue decomposition $A=V\Lambda V^{-1}$. Then,
$\|\exp(At)\|\leq\operatorname*{cond}(V)e^{\lambda_{\min}(A)t}$
@@
$\int_{0}^{t}\|\exp(A\tau)B\|d\tau\leq\frac{\operatorname*{cond}(V)}{|\lambda_{\min}(A)|}\|B\|(1-e^{\lambda_{\min}(A)t}).$
A similar result can be obtained when $A$ is not diagonalizable by using the Jordan form of $A$; see [47, Appendix C.5]. However, since diagonalizable matrices are dense in the space of square matrices and also since $A$ is a design matrix, the diagonalizability assumption is mild.
@@
$\|\tilde{z}(t)\|\leq\|\tilde{z}_{0}\|\operatorname*{cond}(V)e^{\lambda_{\min}(A)t}+\frac{\operatorname*{cond}(V)}{|\lambda_{\min}(A)|}\|B\|(1-e^{\lambda_{\min}(A)t})\big[\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v}\big]$ (32)
@@
Recall $\sigma(t)=h(\bar{x}(t))-h(x(t))-v(t)$ from (29). Then, the solution $\tilde{z}(t;\tilde{z}_{0},\sigma)$ of (29) satisfies
$\|\tilde{z}(t)\|\!\leq\!\|\exp(At)\tilde{z}_{0}\|+\!\int_{0}^{t}\!\|\exp(A\tau)B\sigma(t-\tau)\|\mathrm{d}\tau.$ (33)
@@
$\int_{0}^{t}\|\exp(A\tau)B\sigma(t-\tau)\|\mathrm{d}\tau\hskip 56.9055pt$
$\hskip 56.9055pt\leq\int_{0}^{t}\|\exp(A\tau)B\|\|\sigma(t-\tau)\|\mathrm{d}\tau$
$\hskip 56.9055pt\leq\int_{0}^{t}\|\exp(A\tau)B\|\mathrm{d}\tau\,\max_{\tau\geq 0}\|\sigma(\tau)\|.$ (34)
From (7.1), it follows that $\max_{\tau\geq 0}\|\sigma(\tau)\|\leq\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v}.$ Substituting this in (34) and in (33), and using Lemma 16, we obtain (32). ∎
@@ 7.3 Robustness to the Approximation Error
To analyze the observer’s performance, we employ some measure-theoretic concepts. Recall the general probability measure $\mathcal{D}_{\mathcal{X}}$ over the compact state space $\mathcal{X}$. To describe the asymptotic, long-term statistical distribution of the system’s trajectories, we consider physical invariant measures [21, 53]. For the autonomous nonlinear system (1), $\mu$ is a physical invariant measure if the probability of the state being in any set $X\subseteq\mathcal{X}$ remains constant under the system’s flow [6]. Note that in deterministic dynamical systems, the state evolution itself is not random. However, the randomness originates from the system initialization. Therefore, any probabilistic statement refers to an ensemble of trajectories generated by a distribution of initial conditions.
The primary objective of the observer is to minimize estimation error under the system’s long-term, steady-state operating conditions. For many nonlinear systems, especially the ones we consider in Section 8, the asymptotic behavior is confined to the attractor, whose statistics are governed by the invariant measure $\mu$. The generalization bounds derived in Theorem 14 guarantee that the inverse map error $R_{\mathcal{T}^{*}}(\eta)$ is small only if the distribution $\mathcal{D}_{\mathcal{X}}$ is consistent with the distribution encountered during future operation. If $\mathcal{D}_{\mathcal{X}}$ instead would be some transient initial distribution, minimizing the generalization error would not guarantee accuracy at steady state. In this section, by setting $\mathcal{D}_{\mathcal{X}}=\mu$, we ensure the domain over which we proved the learning guarantee (Theorem 14) is the same domain over which we guarantee stability of the estimation error.
Given the invariant measure $\mu$, the injective KKL transformation map $\mathcal{T}:\mathcal{X}\to\mathcal{Z}$ induces a pushforward measure $\mu_{\mathcal{T}}$ on the observer space $\mathcal{Z}$, defined by $\mu_{\mathcal{T}}(Z)=\mu(\mathcal{T}^{-1}(Z))$, $\forall Z\subseteq\mathcal{Z}$. Because $\mu$ is the invariant measure, the induced measure $\mu_{\mathcal{T}}$ represents the steady-state distribution of the transformed states $z$. The expectation $\mathbb{E}_{z\sim\mu_{\mathcal{T}}}[\cdot]$ with respect to this measure is the formal definition of the steady-state expectation because, for any function $g:\mathcal{Z}\to\mathbb{R}$,
$\mathbb{E}_{z\sim\mu_{\mathcal{T}}}[g(z)]=\int_{\mathcal{Z}}g(z)\mathrm{d}\mu_{\mathcal{T}}(z)=\int_{\mathcal{X}}g(\mathcal{T}(x))\mathrm{d}\mu(x)=\mathbb{E}_{x\sim\mu}[g(\mathcal{T}(x))].$
Since the estimate $\hat{x}(t)$ is obtained from the learned observer (26), we have $x(t)-\hat{x}(t)=\mathcal{T}^{*}(z(t))-\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))$. Thus, the steady-state estimation error is given by
$\xi_{ss}(\zeta)\coloneqq\lim_{t\to\infty}\left[\mathcal{T}^{*}(z(t))-\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))\right]$ (35)
where $\zeta=\lim_{t\to\infty}z(t)$. The high-probability bound on the inverse map error $R_{\mathcal{T}^{*}}$ in Theorem 14 shows that our learning procedure can find an inverse map $\hat{\mathcal{T}}_{\eta}^{*}$ that is, on average, close to the true inverse $\mathcal{T}^{*}$. We now analyze the consequence of this approximation by demonstrating how $R_{\mathcal{T}^{*}}$ directly governs the steady-state performance and robustness of the learned observer (26). We first analyze the noise-free case and then extend the result to uncertain systems.
Theorem 18.
Let Assumptions 2, 6, and 10 hold. Let $\hat{x}(t)$ be the state estimate from the learned KKL observer (26) for the noise-free system (1). Then, the expected squared steady-state estimation error is bounded as follows
$\mathbb{E}_{\zeta\sim\mu_{\mathcal{T}}}\big[\|\xi_{ss}(\zeta)\|^{2}\big]\leq 2R_{\mathcal{T}^{*}}$
where $\xi_{ss}(\zeta)$ is the steady-state error given in (35) and $R_{\mathcal{T}^{*}}$ is the inverse map error (14) bounded by (24).
@@
From (25) and (29), the estimation error $\xi(t)=x(t)-\hat{x}(t)$ satisfies
$\|\xi(t)\|$ $=\|\mathcal{T}^{*}(z(t))-\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))\|$
$\leq\|\hat{\mathcal{T}}_{\eta}^{*}(z(t))-\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))\|+\|\mathcal{T}^{*}(z(t))-\hat{\mathcal{T}}_{\eta}^{*}(z(t))\|$
$\leq\ell_{\eta}\|\tilde{z}(t)\|+\|\mathcal{E}_{\eta}^{*}(z(t))\|$
where $\tilde{z}(t)=\hat{z}(t)-z(t)$ and $\mathcal{E}_{\eta}^{*}(z(t))$ given by (25). Using the inequality $(a+b)^{2}\leq 2a^{2}+2b^{2}$, we have
$\|\xi(t)\|^{2}\leq 2\ell_{\eta}^{2}\|\tilde{z}(t)\|^{2}+2\|\mathcal{E}_{\eta}^{*}(z(t))\|^{2}.$
Since $A$ is Hurwitz, $\tilde{z}(t)$ converges to zero exponentially. Therefore, $2\ell_{\eta}^{2}\|\tilde{z}(t)\|^{2}$ vanishes as $t\to\infty$. By taking the limit on both sides and $\zeta=\lim_{t\to\infty}z(t)$, we obtain
$\|\xi_{ss}(\zeta)\|^{2}\leq 2\|\mathcal{E}_{\eta}^{*}(\zeta)\|^{2}.$
We now take the expectation of this inequality over the steady-state invariant measure $\mu_{\mathcal{T}}$, which gives
$\mathbb{E}_{\zeta\sim\mu_{\mathcal{T}}}[\|\xi_{ss}(\zeta)\|^{2}]\leq 2\mathbb{E}_{\zeta\sim\mu_{\mathcal{T}}}[\|\mathcal{E}_{\eta}^{*}(\zeta)\|^{2}]=2R_{\mathcal{T}^{*}}.$
This concludes the proof. ∎
This theorem provides a link between the learning guarantee in Theorem 14 and the observer’s long-term performance. It shows that the expected steady-state estimation error is directly bounded by the inverse map error $R_{\mathcal{T}^{*}}$. This result is highly practical as Theorem 14 gives a recipe for reducing $R_{\mathcal{T}^{*}}$ (i.e., more data, appropriate network complexity, $\ell_{\eta}$ regularization), whereas Theorem 18 proves that doing so provably improves the average-case performance of the KKL observer’s state estimate.
7.4 Robustness to Uncertainty and Noise
We now analyze the robustness of the learned observer (26) to estimate the state of an uncertain nonlinear system (27). Note that the design method of KKL observers as presented in Sections 3 and 5 remains the same for (27).
Theorem 19.
@@ 7.4 Robustness to Uncertainty and Noise
Let Assumptions 2, 6, 10, and 15 hold. Let $\hat{x}(t)=\hat{\mathcal{T}}_{\eta}^{*}(\hat{z}(t))$ be the estimate from (26) for the uncertain system (27). Suppose $A$ is diagonalizable with eigendecomposition $A=V\Lambda V^{-1}$. Then, the expected squared steady-state estimation error satisfies
$\mathbb{E}_{\zeta\sim\mu_{\mathcal{T}}}\big[\|\xi_{ss}(\zeta)\|^{2}\big]\leq 2R_{\mathcal{T}^{*}}+2\ell_{\eta}^{2}\left(\frac{\operatorname*{cond}(V)}{|\lambda_{\min}(A)|}\|B\|\right)^{2}\left(\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v}\right)^{2}$ (36)
where $\xi_{ss}$ is given in (35), $R_{\mathcal{T}^{*}}$ is bounded by (24), $\overline{w},\overline{v},\psi$ are given in Assumption 15, $\ell_{h}$ is given in (31), $\ell_{\eta}$ is given in (20), and $n_{y}$ is the dimension of the output vector $y(t)$.
@@
We begin from the same squared inequality as in the proof of Theorem 18, given by
$\|\xi(t)\|^{2}\leq 2l_{\eta}^{2}\|\tilde{z}(t)\|^{2}+2\|\mathcal{E}_{\eta}^{*}(z(t))\|^{2}.$
We use Lemma 17 to bound $\|\tilde{z}(t)\|$. As $t\to\infty$, the exponential transient term decays to zero, leaving the steady-state bound
$\limsup_{t\to\infty}\|\tilde{z}(t)\|\leq\frac{\operatorname*{cond}(V)}{|\lambda_{\min}(A)|}\|B\|[\ell_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v}].$
Let $C_{z}=\frac{\operatorname*{cond}(V)}{|\lambda_{\min}(A)|}\|B\|[l_{h}\psi(\overline{w})+\sqrt{n_{y}}\overline{v}]$. Then, the steady-state estimation error is bounded by
$\limsup_{t\to\infty}\|\xi(t)\|^{2}\leq 2l_{\eta}^{2}C_{z}^{2}+\limsup_{t\to\infty}2\|\mathcal{E}_{\eta}^{*}(z(t))\|^{2}.$
Taking the expectation over the steady-state invariant measure yields
$\mathbb{E}_{\zeta\sim\mu_{\mathcal{T}}}[\|\xi_{ss}(\zeta)\|^{2}]\leq 2l_{\eta}^{2}C_{z}^{2}+2\mathbb{E}_{\zeta\sim\mu_{\mathcal{T}}}[\|\mathcal{E}_{\eta}^{*}(\zeta)\|^{2}].$
Substituting $C_{z}$ and $R_{\mathcal{T}^{*}}$, we obtain (36). ∎
This theorem extends our analysis to the uncertain system (27), which includes model uncertainties and sensor noise. It shows that the expected steady-state error is additively composed of two terms: a learning-dependent term and a design-dependent term. The learning-dependent term $R_{\mathcal{T}^{*}}$ is the approximation error for the learned inverse map $\hat{\mathcal{T}}_{\eta}^{*}$, which we can control via the learning process (Theorem 14). The design-dependent term depends on the noise bounds ($\overline{w},\overline{v}$) and the observer’s design parameters ($A,B,\ell_{\eta}$). This term can be minimized by choosing $A$ to be stable ($|\lambda_{\min}(A)|$ large) and well-conditioned ($\operatorname*{cond}(V)$ small [^6] For instance, when $A$ is a diagonal matrix, we have $V=I_{n_{z}}$ and $\operatorname*{cond}(V)=1$.), and by regularizing the learning of $\hat{\mathcal{T}}_{\eta}^{*}$ to keep $\ell_{\eta}$ small. This provides a quantitative link between learning, observer design, and the estimation performance.